OpenStudy (anonymous):

What's the answer too |3x-7|=12

OpenStudy (anonymous):

First, suppose 3x - 7 is greater then 0 and solve 3x - 7 = 12 Then, suppose it is not and solve -(3x - 7) = 12, that is 7 - 3x = 12

OpenStudy (anonymous):

5/3

OpenStudy (anonymous):

You're wrong though. :| @finiteok

OpenStudy (anonymous):

5/3 and 19/3

OpenStudy (anonymous):

@mathway I am not wrong but thank you

OpenStudy (anonymous):

Do what @Andrea_Grossoni said. Solve for x on both equations.

OpenStudy (anonymous):

X is 19/3 and 5/3

OpenStudy (anonymous):

Is what Andrea said the final answer?

OpenStudy (anonymous):

No. Solve for x in: 3x-7=12 and -(3x-7)=12

OpenStudy (anonymous):

@finiteok , the answer 5/3 is wrong, put 5/3 instead of x in the equation and you'll see The answer is -5/3, and of course 19/3

OpenStudy (anonymous):

I wish you were, but you're not. Try again. @finiteok ((but you got one correct though))

OpenStudy (anonymous):

Thanks Andrea!

OpenStudy (anonymous):

She's right! Great job, Andrea! :)

OpenStudy (anonymous):

I hope that "she" was not referred to me. In Italy it is a male name :D Always a pleasure to help, so I don't feel guilty when I get helped!

OpenStudy (anonymous):

My bad. :| I didn't know that Andrea can also be a male's name! Thank you for that information. :D